. What happens in the relativistic addition of velocities formula (derived
in problem (10) when the two velocities are both very small? Try it for
vBA = .01c and vA = .01c. How much does the relativistic result differ
from the Newtonian result (1.2) in this case?
1
Expert's answer
2019-11-11T15:56:35-0500

When the two velocities are both much smaller than the velocity of light, then "v_A v_{BA} / c^2 \ll 1", and this quantity can be neglected compared to unity in the denominator of equation (3.4) for relativistic addition of velocities. In this case, we obtain the Newtonian formula of addition "v_B \approx v_A + v_{BA}". For "v_A = v_{BA} = 0.01 c", the Newtonian formula gives "v_B \approx 0.02 c", while the relativistic formula gives "v_B \approx 0.019998 c", with the difference "\Delta v_B \approx 2 \times 10^{-6} c", or relative difference "\Delta v_B / v_B \approx 2 \times 10^{-6} / 0.02 = 10^{-4}", which is rather small.